OpenStudy (anonymous):

How do I find the values for x and y using m=3 and (1,-4), (x,5), (-2,y)

OpenStudy (mertsj):

\[\frac{5-(-4)}{x-1}=3\]

OpenStudy (mertsj):

\[\frac{9}{x-1}=\frac{3}{1}\]

OpenStudy (mertsj):

\[3x-3=9\]

OpenStudy (mertsj):

\[x=4\]

OpenStudy (mertsj):

Now, can you find y in the same way?

OpenStudy (phi):

the way I would answer this question, is to find the equation of the line. We are given the slope and a point (slope, m=3) and point (1,-4) See http://www.khanacademy.org/video/equation-of-a-line-1?topic=core-algebra once I found the equation, I would plug in numbers to find an x or y. In this case, the equation of the line is y = 3x-7 to find y in the pair (-2,y) replace x with -2 in the equation: y= 3*-2 - 7 to find x in the pair (x,5), replace y with 5: 5=3x-7 and solve for x. If you do not know how to solve an equation see http://www.khanacademy.org/video/simple-equations?topic=core-algebra and following couple of videos